LFBG and LBC Checkout

LFBG notes are on Commissioning wiki

LBC: Fernando and Mauro are here. They brought with them spare parts and a spare controller to work on the Blue tech chip noise problem. Now there are 5 controllers here, and all are interchangeable. If the controller for the Blue Tec chips is faulty, then there is no good spare controller. So far they have not been able to work on the tec chip controller; that is planned for tomorrow.

Today, the LBC Computers were moved from Computer Room A to Computer Room B.

They took a 120 dark this morning and we see no pattern on it (no wedge shape nor 'scallop' along left edge). I took some darks at the start of the night, perhaps before the telescope was on or at least was tracking. I might have seen the wedge pattern but very hard to discern. Not like on the nights before. But when they started to observe with LFBG, I saw the wedge again and the scallop pattern also (see 042146).

Also see 102017, which was a dark taken after the telescope was parked, drives are off.

On new set of darks we see vertical bands of very high, 30K, counts on LBC-Red. These diminish over time. Fernando & Mauro investigating - perhaps these occur when instrument coming out of standby mode (after 10 min of not being used). The bands appear on biases and diminish finally.

-- OlgaKuhn - 17 Sep 2011
  • 360s dark with LBC-Red taken at 05:25 UT:
    r052547.jpg

  • 180s dark with LBC-Red, taken at 10:20 UT after 5 hours idle time:
    r102010.jpg

  • 180s dark with LBC-Red, taken at 10:20 UT after 5 hours idle time:
    r102010.jpg

  • 1st bias after pair of darks, first of which was 102010:
    r103831.jpg

  • 180s dark after set of 25 biases, starting with 103831:
    r110414.jpg

  • 2nd bias after 180s dark 110414...:
    r111200.jpg
I Attachment Action Size Date Who Comment
r052547.jpgjpg r052547.jpg manage 233 K 18 Sep 2011 - 02:20 OlgaKuhn 360s dark with LBC-Red taken at 05:25 UT
r102010.jpgjpg r102010.jpg manage 113 K 18 Sep 2011 - 02:26 OlgaKuhn 180s dark with LBC-Red, taken at 10:20 UT after 5 hours idle time
r103831.jpgjpg r103831.jpg manage 148 K 18 Sep 2011 - 02:28 OlgaKuhn 1st bias after pair of darks, first of which was 102010
r110414.jpgjpg r110414.jpg manage 130 K 18 Sep 2011 - 02:29 OlgaKuhn 180s dark after set of 25 biases, starting with 103831
r111200.jpgjpg r111200.jpg manage 181 K 18 Sep 2011 - 02:31 OlgaKuhn 2nd bias after 180s dark 110414...
Topic revision: r3 - 18 Sep 2011, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback